# An Exploratory Study on Using Large Language Models for Mutation Testing

### v1

@@ 5.1.3. Performance of Using the Same Number of Mutations
In Section 4, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of existing study (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by $\mu$Bert (Degiovanni and Papadakis, 2022), which is 16,785. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 13. Therefore, under fixed quantity conditions, the mutations generated by GPT-3.5-Turbo are still the closest to real bugs in behavior.
@@
Metric GPT-3.5 CL-13b LEAM $\mu$BERT Major PIT
Comp. Rate 62.9% 74.2% 38.9% 18.5% 96.5%
Useless Mut. Ratio 11.0% 39.0% 1.4% 1.4% 0.0%
Real Bug Det. 0.892 0.735 0.675 0.550 0.833 0.437
Coupling Rate 0.133 0.105 0.132 0.132 0.128 0.024
Ochiai Coefficient 0.358 0.178 0.167 0.150 0.341 0.192
@@ 5.2. Implications
First, our study shows that existing LLMs have the potential to generate high-quality mutations for mutation testing with acceptable costs. The greatest advantage of LLMs is their ability to generate mutations that more closely mimic real bugs in behavior. For example, Listing 1 shows two compilable mutations generated by GPT-3.5 in Chart-1, which are theoretically beyond the capabilities of existing approaches. In Mutation-1, the LLM inferences there must exist leftBlock because of the occurrence of rightBlock, and replaces the variable correspondingly. This mutation is not killed by the bug-triggering test. Without natural language understanding capabilities (i.e., understanding the correspondence between “right” and “left”), it is impossible to create such mutations. In Mutation-2, the LLM modifies the else branch to else if, employing an integer variable seriesCount from context to synthesize the condition, showing the capabilities of code structural changing that existing methods cannot achieve. Also, this mutation is alive and can guide test enhancement.
@@ 5.3. Threats to Validity
The selected LLMs, programming language, datasets, and baseline approaches could be a threat to the validity of our results. To mitigate this threat, we adopt the most widely studied models (i.e., GPT and CodeLlama), the most popular language (i.e., Java), and the most popular dataset Defects4J. We also employ state-of-the-art mutation testing approaches as baselines, including learning-based (i.e., $\mu$Bert and LEAM) and rule-based (i.e., PIT and Major).
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014a) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2023b) which includes programs and faults that were made after the release time of the LLMs we use and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. Our results are: GPT, CodeLlama, Major, LEAM, and $\mu$Bert, 282, 77, 67, 386, 39 on the Defects4J dataset while on the ConDefects are 7, 9, 13, 8, 1, respectively, and indicate that on Defects4J GPT and LEAM, approaches tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. $\mu$Bert wields significantly the least number of exact matches, indicating a minimal or no advantage for all these approaches (except from GPT and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interesting, in the ConDefects dataset, which has not been seen by any of the tools, Major has the majority of the exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with $p$-value less than 0.05, indicating their performance is similar on the two datasets.
The different experimental settings may also threaten the validity of our results. To address this threat, we elaborately explore the impacts of prompts, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar.

### v2

@@ 5.1.3. Performance of Using the Same Number of Mutations
In Section 4, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of existing study (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by $\mu$Bert (Degiovanni and Papadakis, 2022), which is 16,785. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 12. Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
@@
Metric GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert Major PIT
Comp. Rate 62.9% 77.4% 70.1% 73.0% 39.3% 74.2% 38.9% 18.5% 96.5% ——
Useless Mut. Rate 11.0% 7.3% 7.0% 8.9% 27.1% 39.0% 1.4% 1.4% 0.0% ——
Real Bug Det. 89.2% 90.8% 91.3% 89.7% 58.0% 73.5% 67.5% 55.0% 83.3% 43.7%
Coupling Rate 13.3% 14.6% 12.6% 13.4% 11.9% 10.5% 13.2% 13.2% 12.8% 2.4%
Ochiai Coefficient 35.8% 45.0% 40.8% 35.3% 15.3% 17.8% 17.0% 15.0% 34.1% 19.2%
5.2. Economic Costs
Closed-source LLMs (e.g., GPT LLMs) are typically billed based on the number of tokens used. As these models continue to evolve, their pricing rates also change, and we here discuss the costs of the proprietary models used in our experiments. For the three models involved, GPT-3.5 incurred a total cost of \$108, GPT-4o cost \$1531, and GPT-4o-Mini cost \$47. Therefore, for every 1,000 mutations, GPT-3.5, GPT-4o, and GPT-4o-Mini cost about \$0.16, \$2.84, and \$0.08, respectively. GPT-4o is significantly more expansive than the others. Given the effectiveness of mutations generated by GPTs, the cost is feasible for large-scale applications.
@@ 5.3. Implications
First, our study shows that existing LLMs have the potential to generate high-quality mutations for mutation testing with acceptable costs. The greatest advantage of LLMs is their ability to generate mutations that more closely mimic real bugs in behavior. For example, Listing 1 shows two compilable mutations generated by GPT-3.5 in Chart-1, which are theoretically beyond the capabilities of existing approaches. In Mutation-1, the LLM inferences there must exist leftBlock because of the occurrence of rightBlock, and replaces the variable correspondingly. This mutation is not killed by the bug-triggering test. Without natural language understanding capabilities (i.e., understanding the correspondence between “right” and “left”), it is difficult to create such mutations. In Mutation-2, the LLM modifies the else branch to else if, employing an integer variable seriesCount from context to synthesize the condition, showing the capabilities of code structural changing that existing methods cannot achieve. Also, this mutation is alive and can guide test enhancement.
@@ 5.4. Threats to Validity
The selected LLMs, programming language, datasets, and baseline approaches could be a threat to the validity of our results. To mitigate this threat, we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek), the most popular language (i.e., Java), and the most popular dataset Defects4J. We also employ state-of-the-art mutation testing approaches as baselines, including learning-based and rule-based.
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014a) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024) which includes programs and faults that were made after the release time of the LLMs we use and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. On Defects4J, the results are GPT-3.5 (469), GPT-4o (290), GPT-4o-Mini (528), DeepSeek (318), StarChat (122), CodeLlama (236), LEAM (590), $\mu$Bert (66), and Major (217), respectively. On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek (32), StarChat (35), CodeLlama (56), LEAM (56), $\mu$Bert (5), and Major (92), respectively. The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. $\mu$Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except from GPTs and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of the exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with $p$-value less than 0.05, indicating their performance is similar on the two datasets.
The different experimental settings may also threaten the validity of our results. To address the threat, we explore the impacts of prompts, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar.
@@ 6. Conclusion
7. Data Availability
All the experimental data, code, and full results (including generated and manually labeled mutations) are publicly available (our, 2024).